A bus makes a trip according to the position–time graph shown in the drawing. What is the average velocity (magnitude and direction) of the bus during each of the segments labeled A, B, and C? Express your answers in km/h.


1
Expert's answer
2021-02-10T01:12:24-0500

a)

"\frac{ 10-40}{1.5 }" = "\frac{30}{1.5}" = 20 km/h (retardation)

b) "\frac{20-10}{2.5-1.5}" = "\frac{10}{1}" = 10 km/hr (acceleration)

c) "\frac{40-20}{3-2.5}" = "\frac{20}{0.5}" = 40 km/h ( acceleration )